OpenStudy (anonymous):

solve using the elimination method. find orderd pair. x-2y=12 4x+2y=8

OpenStudy (callisto):

Add the first equation and the second equation eliminate y. Can you do it?

OpenStudy (anonymous):

x-2y = 12 4x + 2y = 8 5x = 20 then divide both sides by 5 to get x substitute the value of x to any of the two equation for the value of y *kindly see if i'm correct , if im not sorry :)

OpenStudy (anonymous):

let x-2y=12 be equation one. let 4x+2y=8 be equation two. since the numerical coefficients of 'y' are the same, and both of them have opposite signs, we add equation one and two.

OpenStudy (anonymous):